The patient recovery time from a particular surgical procedure is normally distributed with a mean of 5.1 days and a standard deviation of 2.1 days.

What is the probability of spending more than 3 days in recovery? (Round your answer to four decimal places.)
Answer #1

Let, x: Patient recov recovery Hime to Normal (ll -5.1 10 = 2.1) p(x>3) 1-X23) = 1 - Plzc 3-5.1 2.1 = 1 - P(ZL-1) 1 0.1587 Fr